OpenStudy (anonymous):

Write the general solution corresponding to 11pi/6

OpenStudy (anonymous):

The equation \[\cos =\frac{ \sqrt{3} }{ 2 }, 0 \le \theta < 2 \pi \], has solution 11pi/6. Suppose the domain is not restricted. Write the general solution corresponding to 11pi/6.